OpenStudy (mony01):

find a power series representation for the function f(x)=(4+x)/(1-x)

OpenStudy (mony01):

\[f(x)=4+5\sum_{n=1}^{\infty} \] this is also part of the problem

OpenStudy (anonymous):

One way is to know the power series for 1/(1-x) = 1+x+x^2+x^3... and then get a new one by multiplying term by term with (4+x). Alternatively, use Taylor expansion formula with derivatives etc.

OpenStudy (mony01):

i haven't learned the taylor series yet

OpenStudy (anonymous):

Fine. Try my first approach?

OpenStudy (mony01):

can you give me an example to how to start it?